Circular No. 8958

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 FK_{75} (LEMMON-SIDING SPRING) An apparently asteroidal object that was discovered in the course of

the Mt. Lemmon Survey (and given the designation 2008 FK_{75} on MPS 244903; additional observations on MPS 245852) was rediscovered in the course of the Siding Spring survey (again as apparently asteroidal) on July 1 (discovery observations tabulated below) and posted on the Minor Planet Center’s ‘NEOCP’ webpage before the identity was found. J. Young writes that images taken with the Table Mountain 0.61-m f/16 Cassegrain reflector on July 3.2 UT show a 6′′-diameter round, diffuse coma with a slight central condensation; there is no visible tail seen with long or stacked exposures, but the comet was low and in a somewhat-light-polluted area of the sky. Young adds that additional exposures taken with the same telescope by M. Hicks on July 7.2 again show a diffuse, round coma of diameter 6′′ with no tail. R. Holmes (Charleston, IL, U.S.A., 0.81-m f/4.0 astrograph) reports that the object does not appear stellar on his images from July 6.1, the object being somewhat larger than stars of similar brightness).

2008 UT α_{2000} δ_{2000} Mag. Observer
Mar. 31.39920 14^{h}23^{m}35.^{s}16 −18^{o}28′59′′.3 19.4 Mt. Lemmon
July 1.43226 13 56 58.34 −11 44 19.0 18.9 Siding Spring

Additional astrometry and the following parabolic orbital elements by G. V. Williams appear on MPEC 2008-N18.

T = 2010 Sept.30.8879 TT  ω = 80.^{o}3140
Ω = 218.2486 }2000.0
q = 4.524280 AU  i = 61.2159

CATALOGUE OF COMETARY ORBITS 2008 A new 195-page edition — the seventeenth — of the MPC/CBAT Cat-

alogue of Cometary Orbits has now been published. The catalogue includes 3815 orbits for 3708 cometary apparitions, and there are tabulations of the osculating elements for six standard epochs beginning 2007 Oct. 27.0 for periodic comets. The 195-page catalogue sells for $60.00 ($80.00 for airmail delivery outside North America). There is also an e-mail edition (with just the basic orbital information) for $80.00. Orders should be sent to the address in the header of this Circular.

2008 July 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green